Its a word problem
A person holds a pistol straight upwards and fires. The initial velocity of must bullets is around 1200 ft/sec. The height of the bullet is a function of time and is described by h(x)=_=-16t^2+1200t. How long,, after the gun is fired, does the person have to get out of the way of the bullet falling from the sky?

Answer by Alan3354(69443)

h(x) = -16t^2+1200t
Find when h = zero.
-16t^2+1200t = 0
t = 0 (when fired)
t = 1200/16 = 75 seconds (bullet hits the ground)